# Evaluations of Machine Learning Privacy Defenses are Misleading

### v1

@@ 1. Introduction
    Current membership inference evaluations fail to reflect a model’s privacy on the most vulnerable data, and instead aggregate the attack success over a population. But privacy is not an average-case metric! (Steinke and Ullman, 2020) We show that a blatantly non-private defense that fully leaks one training sample passes existing evaluations (even with recent proposals to report an attack’s true-positive rate at low false-positive rates (Carlini et al., 2022a)).
@@
Our work adds to the growing literature on pitfalls in evaluations of ML privacy defenses (Choquette-Choo et al., 2021; Tramer et al., 2022; Carlini et al., 2022b). We aim to provide researchers a more principled evaluation framework, and an overview of pitfalls and misconceptions in existing evaluations. To promote reproducible research, we release all code for our evaluation methodology and our implementation of each empirical defense we study.111 https://github.com/ethz-spylab/misleading-privacy-evals
@@ 4.1. Focus on the Most Vulnerable Samples
Thus, in existing evaluations, this defense can be made arbitrarily private by increasing the size of $D$. Evaluating on more samples indeed improves the defense’s privacy on average across the population (i.e., a smaller proportion of the dataset is at risk of privacy leakage); yet, from the individual perspective of $\hat{x}$, the defense never provides any privacy at all.
@@
$\text{TPR}({T})\coloneq\Pr_{x\sim D_{\textrm{in}}}{[\mathcal{A}(f,x)>{T}]},\quad\text{FPR}({T})\coloneq\Pr_{{x\sim D_{\textrm{out}}}}{[\mathcal{A}(f,x)>{T}]},$
@@
If we train $S$ models and evaluate the attack on $C$ samples, we thus want $S\cdot C\gg 1/\alpha$, introducing a tradeoff between tightness of the privacy bound and computational efficiency. Existing works mainly focus on two extremes:
@@
Our (illustrative) approach in Figure 2(a) follows the latter extreme: we used 20,000 models to tightly approximate the per-sample attack success at low FPRs. Such an approach is of course impractical in general, especially since many privacy defenses add computational overhead over standard training.
We thus adopt a natural middle ground: we evaluate membership inference on a small set of $1\ll C\ll|D|$ samples (called “canaries” (Carlini et al., 2019)), where each canary is inserted independently at random in the training data of a small number $S$ of models. The evaluation then reports membership inference success only over the canary set, ignoring the remaining data.
@@ 5.1. Defenses
Self-Supervised Learning is a technique to learn feature representations from unlabeled data. Given a labeled dataset $\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$, the SSL defense first trains a feature encoder $\phi$ in an unsupervised fashion, using only the features $\{x_{1},\dots,x_{n}\}$. We consider two popular methods, SimCLR (Chen et al., 2020) and MoCo (He et al., 2020), both employing a contrastive loss which ensures that different augmentations of an input yield similar features (i.e., $\phi(x)\approx\phi(\texttt{aug}(x))$). Then, in a second stage, we train a linear classifier $f(x)=W\phi(x)+b$ on top of the frozen encoder, using the full labeled training set and a standard cross-entropy loss. SSL is not explicitly a defense against membership inference, but has received a lot of study regarding its privacy (Liu et al., 2021; He and Zhang, 2021; Ko et al., 2023). We include this SSL-based defense to illustrate a shortcoming of a naive privacy evaluation that applies a LiRA-like attack out-of-the-box without accounting for the unsupervised nature of the encoder’s training.
@@ 5.5. Strong DP-SGD Baselines
We tune the hyperparameters of the medium utility baseline in the same way as for all case studies (see Section 5.2), that is, to maximize privacy (measured by $\epsilon$ at $\delta=10^{-5}$) subject to at least $88\%$ CIFAR-10 test accuracy. Crucially, we do not enforce meaningful DP guarantees; the provable privacy guarantees $\epsilon$ of all our DP-SGD baselines are in the thousands. For the high utility baseline, we rely on recent scaling laws (Sander et al., 2023) to increase the medium baseline’s utility at the cost of privacy (primarily by decreasing batch size while carefully scaling noise). See Table 3 for all hyperparameters, and the average CIFAR-10 test accuracy over 64 models.
@@ A.3. Validation of Our Auditing Setup
Figure 8. Our threat model yields comparable privacy estimates to prior work. Recall that prior work trains shadow models by randomly sampling the membership of each dataset point. In contrast, we only sample the images that we audit, and keep the remaining parts of the training set fixed. While this gives the attacker more power than in prior evaluations (i.e., the attacker knows a large fraction of the target model’s training set), we show this has a negligible effect on evaluation outcomes. For audit sets of size either 250 or 500, we compare the setting where we resample each image (solid lines) or where we only resample the audit set (dashed lines) and keep $\approx 25{,}000$ images fixed (to ensure training sets of similar sizes in both settings). Our audit setting (dotted line) just samples the audit set, and always uses the remaining CIFAR-10 training data (49,500 images) to train all shadow models. Our approach never overestimates privacy leakage.
@@ A.5. Canaries
Table 2. Canaries have to be adapted to the defense. Mislabeled samples are powerful canaries for most defenses that rely on a standard supervised learning paradigm, but fail to reliably mimic the leakage of the most vulnerable samples for SELENA or SSL.
@@ B.1. Investigating Adaptive Attacks and OOD Effectiveness for SSL
Figure 10(a) shows the performance of our adaptive attack for MoCo on a population level (analogous to the SimCLR results in Figure 4(a)). We again find that using our adaptive score $\operatorname{\mathcal{S}}(f;x)$ in LiRA mildly increases the TPR@0.1% FPR over standard confidence-based scores in a black-box setting (2.0% to 3.6%), and significantly in a white-box setting (to 23.6%, more than an $11x$ increase).

### v2

@@ Abstract.
machine learning; privacy; audit; membership inference; DP-SGD
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690194isbn: 979-8-4007-0636-3/24/10ccs: Computing methodologies Machine learningccs: General and reference Evaluationccs: Security and privacy
@@ 1. Introduction
    Current membership inference evaluations fail to reflect a model’s privacy on the most vulnerable data, and instead aggregate the attack success over a population. But privacy is not an average-case metric! (Steinke and Ullman, 2020) We show that a blatantly non-private defense that fully leaks one training sample passes existing evaluations (even with recent proposals to report an attack’s true positive rate at low false positive rates (Carlini et al., 2022a)).
@@
Our work adds to the growing literature on pitfalls in evaluations of ML privacy defenses (Choquette-Choo et al., 2021; Tramer et al., 2022; Carlini et al., 2022b; Kaplan et al., 2024; Li et al., 2024). We aim to provide a more principled evaluation framework, and an overview of pitfalls and misconceptions in existing evaluations. To promote reproducible research, we release all code for our evaluation methodology and our implementation of each empirical defense we study.111 https://github.com/ethz-spylab/misleading-privacy-evals
@@ 4.1. Focus on the Most Vulnerable Samples
Thus, in existing evaluations, this defense can be made arbitrarily private by increasing the size of $D$. Evaluating on more samples indeed improves the defense’s privacy on average across the population (i.e., a smaller proportion of the dataset risks privacy leakage); however, from the individual perspective of $\hat{x}$, the defense never provides any privacy at all.
@@
$\text{TPR}({T})\coloneq\,\,\Pr_{\mathclap{\begin{subarray}{c}x\sim D_{\textrm{in}}\\ f\sim\text{Train}(D_{\textrm{in}})\end{subarray}}}{\,[\mathcal{A}(f,x)>{T}]},\quad\text{FPR}({T})\coloneq\,\,\Pr_{\mathclap{\begin{subarray}{c}{x\sim D_{\textrm{out}}}\\ f\sim\text{Train}(D_{\textrm{in}})\end{subarray}}}{\,[\mathcal{A}(f,x)>{T}]},$
@@
If we train $S$ models and evaluate the attack on $C$ samples, we thus want $S\cdot C\gg 1/\alpha$. This introduces a tradeoff between the tightness of the privacy bound and computational efficiency.
Existing works mainly focus on two extremes:
@@
Our (illustrative) approach in Figure 2(a) follows the latter extreme: we used 20,000 models to tightly approximate the per-sample attack success at low FPRs. This approach is generally impractical, especially since many privacy defenses add computational overhead.
We thus adopt a natural middle ground: we evaluate membership inference on a small set of $1\ll C\ll|D|$ samples (called “canaries” (Carlini et al., 2019)), where each canary is inserted independently at random in the training data of a small number $S$ of models. The evaluation then reports membership inference success only over the canary set, ignoring the remaining data. Our approach resembles recent efforts to parallelize DP auditing (Pillutla et al., 2023).
@@ 5.1. Defenses
Self-Supervised Learning is a technique to learn feature representations from unlabeled data. Given a labeled dataset $\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$, the SSL defense first trains a feature encoder $\phi$ in an unsupervised fashion, using only the features $\{x_{1},\dots,x_{n}\}$. We consider two popular methods, SimCLR (Chen et al., 2020) and MoCo (He et al., 2020), both employing a contrastive loss which ensures that different augmentations of an input yield similar features (i.e., $\phi(x)\approx\phi(\texttt{aug}(x))$). Then, in a second stage, we train a linear classifier $f(x)=W\phi(x)+b$ on top of the frozen encoder, using the full labeled training set and a standard cross-entropy loss. SSL is not explicitly a defense against membership inference, but has received a lot of study regarding its privacy (Liu et al., 2021; He and Zhang, 2021; Ko et al., 2023). We include this SSL-based defense to illustrate a shortcoming of a naive privacy evaluation that applies a LiRA-like attack out-of-the-box without accounting for the unsupervised nature of the encoder.
@@ 5.5. Strong DP-SGD Baselines
We tune the hyperparameters of the medium utility baseline in the same way as for all case studies (see Section 5.2), that is, to maximize privacy (measured by $\epsilon$ at $\delta=10^{-5}$) subject to at least $88\%$ CIFAR-10 test accuracy. Crucially, we do not enforce meaningful DP guarantees; the provable privacy guarantees $\epsilon$ of all our DP-SGD baselines are in the thousands. For the high utility baseline, we rely on recent scaling laws (Sander et al., 2023) to increase the medium baseline’s utility at the cost of privacy (primarily by decreasing batch size while carefully scaling noise). See Table 3 for all hyperparameters and the average CIFAR-10 test accuracy over 64 models.
@@ A.3. Validation of Our Auditing Setup
Figure 8. Our threat model yields comparable privacy estimates to prior work. Recall that prior work trains shadow models by randomly sampling the membership of each dataset point. In contrast, we only sample the images that we audit, and keep the remaining parts of the training set fixed. While this gives the attacker more power than in prior evaluations, we show this has a negligible effect on evaluation outcomes. For audit sets of size either 250 or 500, we compare the setting where we resample each image (solid lines), or where we only resample the audit set and keep $\approx 25{,}000$ images fixed (dashed lines), and our audit setting (dotted line) where we just sample the audit set, and always include the remaining CIFAR-10 training data (49,500 images). Our approach never overestimates privacy leakage.
@@ A.5. Canaries
Table 2. Canaries have to be adapted to the defense. Mislabeled samples are strong canaries for defenses that rely on a standard supervised learning paradigm, but fail to mimic leakage of the most vulnerable samples for SELENA or SSL.
@@ B.1. Investigating Adaptive Attacks and OOD Effectiveness for SSL
Figure 10(a) shows the performance of our adaptive attack for MoCo on a population level (analogous to the SimCLR results in Figure 4(a)). We again find that using our adaptive score $\operatorname{\mathcal{S}}(f;x)$ in LiRA mildly increases the TPR@0.1% FPR over standard confidence-based scores in a black-box setting (2.0% to 3.6%), and significantly in a white-box setting (to 23.6%, more than an $\times1$ increase).